Vickey has just been fouled and is in a one-and-one free throw situation. This means that she will try one shot. If she makes the shot, she gets to try a second shot. If she misses the first shot, she does not get to try a second shot. Vickey's free throw percentages is 70%. Which of the following do you think is most likely to happen?
a) She will score zero points (she misses both shots)
b) she will score one point(make the 1st and miss the 2nd shot)
c) she will score two points
Please help! I am struggling, I don't understand it. Any help, I would appreciate it! Thank you.

Answer by stanbon(75887)

Vickey's free throw percentages is 70%. Which of the following do you think is most likely to happen?
a) She will score zero points (she misses both shots)
b) she will score one point(make the 1st and miss the 2nd shot)
c) she will score two points
---------------
P(she makes both shots) = 0.7^2 = 0.49
P(she misses both shots) = 0.3^2 = 0.09
P(she makes the 1st shot and misses the 2nd) = 0.7*0.3 = 0.21
---------------------
The highest probability is that she will make both shots.
Ans: c
==================
Cheers,
Stan H.